OpenStudy (firejay5):

9. An object is fired straight up from the top of a 200-foot tower at a velocity of 80 feet per second. The height, h(t) of the object t seconds after firing is given by h(t) = -16t^2 + 80t + 200. A. Find the maximum height reached by the object and the time that the height is reached. B. Interpret the meaning of the y-intercept in the context of this problem.

OpenStudy (anonymous):

If you were to graph it on your calculator (or wolfram) the answer to A. is the y value at the maximum point.

OpenStudy (anonymous):

y-intercept you have to think about it. that is at time 0. Right?! Because the y-intercept is when x=0. So at time 0 what is the height? Hint: it is the initial height

OpenStudy (firejay5):

Is it 299.95473???? ????

OpenStudy (anonymous):

idk I didn't do it

OpenStudy (firejay5):

that number is y, but I don't think it's the y-intercept

OpenStudy (anonymous):

\[h(t) = -16t^2 + 80t + 200\]\[h'(t) = -32t + 80\]\[0 = -32t + 80\]\[32t=80\]\[t=2.5\]\[h(t) = -16(2.5)^2 + 80(2.5) + 200\]\[h(t)=300\]

OpenStudy (anonymous):

300 is the answer to A

OpenStudy (anonymous):

reread my initial post for B

OpenStudy (firejay5):

It says it need the height and the time

OpenStudy (anonymous):

hate to say it but 300 is an exact answer. I don't know how wolfram got its 299.whatever The time is the x value.

OpenStudy (firejay5):

wolframalpha got 300; t = \[\frac{ 5 }{ 2 }\]

OpenStudy (anonymous):

which is 2.5, which is what I got. See how quick calculus is?!

OpenStudy (anonymous):

I want to make sure you are understanding. Do you know why time is the x-value and height is the y value?

OpenStudy (firejay5):

explain it real quick

OpenStudy (anonymous):

\[h(t) = -16t^2 + 80t + 200\]h stands for height in this problem. and h depends on t. That's what h(t) means. You say it as "h of t". Means height changes as t changes. t stands for time. say we had our finger on the x-axis and moved it to the right. we are increasing the time which is increasing the height on the y-axis. did this help

OpenStudy (anonymous):

I should have probably said changing the height as time changes, because the height goes up then back down

OpenStudy (firejay5):

okay I get it now partially

OpenStudy (anonymous):

just remember that h(t) or f(x) or anything of anything is usually always referring to the y-axis. and the of something or t or x is the x-axis

OpenStudy (firejay5):

I am pretty sure we'll get more in to that later in the year maybe

OpenStudy (anonymous):

definitely

OpenStudy (anonymous):

Good luck

OpenStudy (firejay5):

are you sure we'll learn it in Algebra2

OpenStudy (anonymous):

definitely

OpenStudy (firejay5):

What was your initial post for B

OpenStudy (anonymous):

the initial height is the height of the beginning. That is at time zero

OpenStudy (anonymous):

building not beginning

OpenStudy (firejay5):

A is 300 - maximum height and 5/2 - t(time) B was what???

OpenStudy (anonymous):

height of the building.

OpenStudy (firejay5):

so is B the height of the building

OpenStudy (firejay5):

You know what @ChmE

OpenStudy (firejay5):

Thanks a whole lot man I hope you can help me out later sometime! :D